Two parents come to see you for genetic counseling. Dad is sickle cell carrier (sickle cell is a recessive genetic disease) and has a normal cholesterol. Mom is also a carrier for sickle cell and has struggled with high cholesterol, and a test by her doctor recently showed that she is heterozygous for hypercholesterolemia (hypercholesterolemia is a genetic disease that shows incomplete dominance). What are the odds of them producing a child with sickle cell disease?-------A child who is a carrier for the sickle cell trait?--------A child with heterozygous hypercholesterolemia?-------A child with homozygous hypercholesterolemia?----------What are the odds of producing a child heterozygous for both?--------
Answer #1

Let the normal allele be "S" and sickle cell allele be "s" . Also, normal allele for cholesterol be C1 and hypercholesterolemia allele be C2.

Note that S is dominant over s allele, but C1 and C2 alleles show incomplete dominance, which means that heterozygous condition has an intermediate phenotype.

Dad

(sickle cell carrier, normal cholesterol)

Genotype of Dad : SsC1C1

Mom

(sickle cell carrier, heterozygous for (hypercholesterolemia)

Genotype of Mom: SsC1C2

Thus, the required cross is SsC1C1 × SsC1C2.

Let us set up punnet squares for each gene as follows:

Ss × Ss

Cross for

Sickle cell

S s
S SS (normal) Ss (carrier)
s Ss (carrier) ss (affected)

For cholesterol gene, C1C1 × C1C2

C1 C2
C1 C1C1 (normal) C1C2 (heterozygous forhypercholesterolemia)

Ans (a) :

The odds of producing a sickle cell child (ss) is 1/4 or 0.25.

Ans (b) :

Probability of child who is carrier for sickle cell trait (Ss) is 1/2 or 0.5.

Ans (c) :

Probability of a child with heterozygous hypercholesterolemia (C1C2) is 1/2 or 0.5.

Ans (d) :

Probability of a child with homozygous hypercholesterolemia (C2C2) is zero.

Ans (e) :

The odds of producing a child heterozygous for both (SsC1C2) can be determined by multiplying the individual probability of heterozygous condition for each trait.

= p(Ss) × p(C1C2)

=(1/2) × (1/2)

= 1/4 or 0.25.

​​​​​

​​​​

​​​​​​